OpenStudy (anonymous):

solve by substitution 2x-3y=-1 y=x-1

OpenStudy (anonymous):

2x-3(x-1)=-1 2x-3x+3=-1 x=4 y=x-1 y=4-1 y=3 (4,3)

OpenStudy (anonymous):

2x - 3(x-1) = -1 2x - 3x +3 = -1 -x = -1 -3 -x = -4 x = 4 y = 4-1 y = 3